help me solve this, please. 7x+x(x-5)=0. I'm not sure what to put in place of the x
Answer by nerdybill(7384)

7x+x(x-5)=0
7x+x^2-5x=0
x^2 + 2x = 0
x(x+2) = 0
Setting each term on the left to zero we get:
x = {0, -2}